# Teager-Kaiser Energy Methods for EEG Feature Extraction in Biomedical Applications

### v1

@@ 3.2 Preprocessing
Table 2: 5-fold Mean Classification Performance and Standard Deviation on BCI-IV 2a.
@@
range of each canonical band we applied filterbanks of 25 filters (SEED, BCI-IV 2a) and 12 filters (TUEP, due to computational constraints), see also Table 1.
3.3 Baseline Feature Extraction
To enable a comparative evaluation of the proposed approach and to assess its potential complementarity with established EEG descriptors, we re-implement a standard feature extraction pipeline for EEG signals, wherein we first isolate the desired frequency bands with a 10th-order Butterworth filter, and then extract a set of baseline features from the isolated bands (or the raw signal). These features include the Power Spectral Density (PSD) and the Signal Energy of the signal. PSD is computed using the Welch method, yielding $f_{s}/2$ features per signal, where $f_{s}$ denotes the sampling frequency. The Signal Energy, defined as the squared magnitude of the raw signal at any point in time and commonly used in signal processing as a direct measure of instantaneous energy, is employed here. Similar to the proposed TKEO statistical descriptors, the mean Signal Energy and the mean Relative Signal Energy features are extracted.
For clarity, we define the term Combined features as the set consisting of the TKEO features, the PSD, mean Signal Energy, and mean Relative Signal Energy, concatenated.
3.4 Evaluation Protocol
We evaluate the utility of TKEO-based features on three classification tasks: motor imagery (BCI-IV 2a dataset), emotion recognition (SEED dataset) and epilepsy detection (TUEP dataset). A Subject Independent setting is presented, where models are trained on a subset of subjects and tested on unseen individuals. This setting reduces the risk of confounding due to subject- or recording-specific biases and better assesses cross-subject generalization. While we also conducted experiments in a Subject Dependent setting (i.e., training and testing on different trials from the same subject), we report only the independent results here, as they better reflect real-world applications and due to space constraints.
Features from all EEG channels were concatenated and the resulting feature vectors were standardized with a Standard Scaler and classified using a Random Forest classifier with 100 estimators, chosen for computational efficiency. We deliberately employed Random Forest, as it is a standard classifier that allows us to assess the discriminative power of the extracted features. This choice ensures that classification performance reflects the quality of the proposed TKEO-based features, rather than the capacity of complex machine learning models. Evaluation is performed with stratified 5-fold cross-validation. SEED and BCI-IV 2a are balanced, whereas TUEP is imbalanced, with 77.7% epileptic recordings; to mitigate this, we report metrics robust to class imbalance, namely Balanced Accuracy and ROC-AUC.
4 Results & Discussion
Table 3: 5-fold Mean Classification Performance and Standard Deviation on SEED.
@@ 4 Results & Discussion
Table 4: 5-fold Mean Classification Performance and Standard Deviation on TUEP.
@@
4.1 Comparison to Baselines
The classification results for all features are summarized in Tables 24. We report results using features extracted directly from the raw signal (raw-band) as well as from the fusion of features computed across all canonical frequency bands (fused-band). TKEO features generally outperform the baseline features (PSD and Signal Energy) in motor imagery and epilepsy detection tasks, whereas in emotion recognition, the baselines achieve higher performance.
As shown in Table 2, TKEO features consistently outperform all other feature sets in the motor imagery task, with the fused-band achieving the best overall performance (39.8% Balanced Accuracy and 66.2% ROC-AUC), substantially exceeding chance level. Moreover, integrating TKEO features with baselines yields performance gains over the baselines alone in most cases, highlighting their complementary nature. In the SEED dataset (Table 3), baselines generally outperform TKEO features, except for the raw-band ROC-AUC (82.0%) of Combined set, reflecting strong class separability. In the TUEP dataset (Table 4), Combined features consistently achieve the highest performance. In particular, the fused-band configuration achieves the highest Balanced Accuracy (76.2%)—approximately 15% higher than other feature sets—and the highest ROC-AUC (82.2%). TKEO features alone attain performance comparable to the baselines, further highlighting their complementary contribution when combined.
Overall, TKEO features show stronger relative performance in motor imagery and epilepsy detection compared to emotion recognition, underscoring their suitability for tasks where transient temporal changes are important and their limitations when the discriminative information is not as temporally localized. Additionally, the fused-band consistently outperforms the raw-band within TKEO features, reflecting the advantage of aggregating information across frequency bands while preserving the narrowband characteristics of TKEO.
4.2 Ablation Study
In our study, EEG signals were decomposed into frequency bands using a filterbank with a fixed number of filters. To examine the impact of the number of filters, we performed an ablation study on the BCI-IV 2a dataset, repeating the analysis with varying filter counts. The results in Table 5 show that the 50-filter configuration achieves the overall highest performance, yielding a Balanced Accuracy of 41.0% and an ROC-AUC of 66.8% in fused-band. This suggests that a finer filterbank provides higher spectral resolution, enabling the TKEO to capture narrowband EEG dynamics that are otherwise averaged out with fewer filters. To further expand our study, we investigate the impact of each frequency band to the fused-band overall accuracies. As shown in Fig. 2, the SHAP [shap] heatmap highlights elevated contributions in sensorimotor regions (C3, C4, CP3, CP4) in Alpha–Beta bands that capture event-related desynchronization/synchronization (ERD/ERS) during imagined movements. This pattern aligns with well-established neurophysiological findings of motor imagery [scherer2018].
Table 5: 5-fold Mean Classification Performance and Standard Deviation on BCI-IV 2a of TKEO Feature for different Filterbanks.
# Filters Balanced Accuracy (%) ROC-AUC ($\mathbf{\times 10^{2}}$)
Raw Fused Raw Fused
@@ 4.2 Ablation Study
Fig. 2: Absolute SHAP values of frequency bands indicating their contributions to fused-band TKEO predictions across EEG channels in BCI-IV 2a; higher SHAP values represent greater contribution.
5 Conclusion
In this paper, we investigated the utility of the Teager-Kaiser Energy Operator on three EEG classification tasks, namely motor imagery, emotion recognition, and epilepsy detection. A Gabor filterbank was applied within each canonical band to extract band-specific signals, followed by the application of TKEO to extract statistical features, and instantaneous modulation features using the DESA-1 demodulation algorithm. Experimental results demonstrate that TKEO-based features achieve competitive performance compared to baseline measures, improving accuracy in motor imagery and epilepsy detection tasks, while performing on par with energy-based features on emotion recognition. Feature fusion across frequency bands further improves performance, highlighting the narrowband sensitivity of TKEO. Overall, our analysis indicates that TKEO is a valuable tool for EEG signal analysis, both as a standalone method and as a complementary feature extraction technique. A promising direction for future work is the integration of TKEO into deep learning (DL) frameworks, combining the predictive power of DL with physiologically meaningful, interpretable representations.

### v2

@@ IV-A Main Results
TABLE I: 5-fold Mean Classification Performance and Standard Deviation on BCI-IV 2a (Motor Imagery).
@@
[…]
TABLE II: 5-fold Mean Classification Performance and Standard Deviation on SEED (Emotion Recognition).
@@
[…]
TABLE III: 5-fold Mean Classification Performance and Standard Deviation on TUEP (Epilepsy Detection).
@@
[…]
IV-B Additional Experiments
In our study, EEG signals were decomposed into frequency bands using a filterbank with a fixed number of filters. To examine the impact of filter count, we performed an ablation study on the BCI-IV 2a dataset. As shown in Fig. 2, performance exhibited minor variations across configurations, with 12 filters achieving the highest accuracy for the raw-band and 50 for the fused-band. To further expand our study, we investigated the impact of each frequency band to the fused-band overall accuracies. As shown in Fig. 3, the SHAP [23] heatmap highlights elevated contributions in sensorimotor regions (C3, C4, CP3, CP4) in Alpha-Beta bands that capture event-related desynchronization/synchronization (ERD/ERS) during imagined movements. This pattern aligns with well-established neurophysiological findings of motor imagery [24]. We also evaluated the main pipeline restricted to sensorimotor regions (C3, Cz, C4, CP3, CPz, CP4) across different bands. The Alpha-Beta combination achieved the highest performance for this subset (40.9% Balanced Accuracy, 67.4% ROC-AUC; Table IV), indicating that these channels and bands retain discriminative information while reducing feature dimensionality.
Figure 2: Performance of different filterbanks on BCI-IV 2a.
Figure 3: Absolute SHAP values of frequency bands indicating their contributions to fused-band TKEO predictions across EEG channels in BCI-IV 2a; higher SHAP values represent greater contribution.
TABLE IV: Performance for Selected Channels on BCI-IV 2a.
Band Balanced Accuracy (%) ROC-AUC ($\mathbf{\times 10^{2}}$)
[…]
V Conclusion
In this paper, we investigated the utility of the Teager-Kaiser Energy Operator on three EEG classification tasks, namely motor imagery, emotion recognition, and epilepsy detection. A Gabor filterbank was applied within each canonical band to extract band-specific signals, followed by the application of TKEO to extract statistical features, and instantaneous modulation features using the DESA-1 demodulation algorithm. Experimental results demonstrate that TKEO-based features achieve competitive performance compared to baseline measures and contribute complementary information, improving accuracy in motor imagery and epilepsy detection tasks, while performing on par with energy-based features on emotion recognition. Feature fusion across frequency bands further improves performance, highlighting the narrowband sensitivity of TKEO. Overall, our analysis indicates that TKEO is a valuable tool for EEG signal analysis, both as a standalone method and as a complementary feature extraction technique. A promising direction for future work is the integration of TKEO into deep learning (DL) frameworks, combining the predictive power of DL with physiologically meaningful, interpretable representations.